OpenStudy (anonymous):

y/y+7=2 How do I know what y is? please explain, thx

OpenStudy (zzr0ck3r):

\(\frac{y}{y+7}=2\) multiply both sides by y+7 \(y=2(y+7)=2y+14\) subtract y from both sides, subtract 14 from both sides \(y=\large ?\)

OpenStudy (anonymous):

Thanks!